# Conditional Feature Importance revisited: Double Robustness, Efficiency and Inference

### v4

@@ 5 Experiments
To compare importance estimators, we work in a regression setting with the standard quadratic loss. We study both linear and more complex settings. In these cases, it is possible to compute explicitly TSI (see Appendix K). Figures 1 and 2 are run over 50 and 100 repetitions respectively. The code is available at https://anonymous.4open.science/r/Sobol-CPI-D9CB.
We compare the proposed method (with varying $n_{\mathrm{cal}}$ values) to LOCO-W—the data-splitting version from Williamson et al., (2023)—and LOCO-HD from Verdinelli and Wasserman, (2024), which, to the best of our knowledge, are the only methods aligned with the same interpretability goals. Our primary focus is on the bias in estimating important and non-important covariates (i.e., the double robustness property), and on the ability to perform powerful, type-I-error-controlled variable selection. For a broader discussion of how $n_{\mathrm{cal}}$ balances nonparametric efficiency (variable importance) and double robustness (variable selection), we refer to Section J.2.
@@
we study a nonlinear setting similar to the one of Bénard et al., (2022): $y=X_{0}X_{1}\mathbb{I}_{X_{2}>0}+2X_{3}X_{4}\mathbb{I}_{X_{2}<0}$, with $X\sim\mathcal{N}(\mu,\Sigma)$, where $\Sigma_{i,j}=0.6^{|i-j|}$, $p=50$, and $\mu=\mathbf{0}$. We use a simple Lasso for $\widehat{\nu}_{-j}$ and a Super-Learner van der Laan et al., (2007) comprising Random Forests, Lasso, Gradient Boosting, and SVM for $\widehat{m}$ and $\widehat{m}_{-j}$. For computational reasons, in the data-splitting version (LOCO-W), we used only Gradient Boosting. We applied Sobol-CPI with $n_{\mathrm{cal}}=1$ and $n_{\mathrm{cal}}=100$.
In Figure 1, we observe that even if we use a complex model to estimate $\widehat{m}_{-j}$, but a simple one to estimate $\widehat{\nu}_{-j}$, we not only achieve a more computationally efficient estimate but also obtain better results with Sobol-CPI than with LOCO. First, for an important covariate (illustrated here by $X_{0}$), Sobol-CPI and LOCO-HD yield similar results, but Sobol-CPI without refitting all the super-learners. Furthermore, for null covariates (such as $X_{6}$), Sobol-CPI shows no bias. Then, we see that Sobol-CPI performs better in identifying important covariates. Finally, the double robustness property of Sobol-CPI is evident, as with $n_{\mathrm{cal}}=100$, it still does not assign importance to the null covariates.
Figure 1: Double robustness for complex learners: left to right: TSI estimates for an important covariate ($X_{0}$) and a null covariate ($X_{6}$); AUC for an importance-based variable selection; bias for null covariates. Sobol-CPI converges at similar rates to TSI and, under the null, it converges faster.
@@
As discussed in Section 4.3, to test the null hypothesis for each covariate, a correction is necessary to address the variance decrease under the null hypothesis. In Figure 2, we present the power of the test using bootstrap variance estimation with a linear correction term. Other approaches, along with analyses of type-I error, computation time, and additional settings (e.g., different correlations and polynomial settings), are discussed in Section J.4.
First, LOCO-W lacks power and it requires a large sample size to control the type-I error. We find that Sobol-CPI also reduces the bias for the non-null covariates. Furthermore, it leverages its double robustness property to achieve the largest power. Finally, the type-I error is controlled at the target level $0.05$ accross all the procedures but the data-splitting version of Williamson et al., (2023).
Figure 2: Statistical Inference on variable importance in a linear setting with correlation: AUC for variable selection accuracy, bias in non-null TSI estimation, power and type-I error. Sobol-CPI(1) provides the most powerful test. Using the corrected variance, the type-I error is controlled.
@@ 6 Conclusion
In this article, we revisited Conditional Feature Importance (CFI). First, we considered the estimation of the conditional distribution, which had been completely ignored in the literature. We then studied both the null and alternative regimes. To explain the good performance of CFI in correctly identifying null features, we proposed a double robustness property. This property means that the method benefits from both the conditional sampler and the model: asymptotically, the sampler draws from the input distribution, so the two sources of error compensate as they share the same distribution, while the model generally induces implicit variable selection during optimization.
For the alternative regime, we modified the CFI and introduced the Sobol-CFI, which we proved to be nonparametrically efficient. We also addressed the bias by proposing a correction that establishes the link between LOCO (refitting) and CFI (perturbation). Finally, we showed how to perform valid inference within the procedure. Overall, we provided a detailed and comprehensive account of the behavior of CFI.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes,https://anonymous.4open.science/r/Sobol-CPI-D9CB]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Appendix
@@ Appendix C Conditional sampling proofs
Also, we have that $X^{j}|\left(X^{-j}=x^{-j}\right)=\epsilon_{j}+\nu_{-j}(x^{-j})\overset{{\mathrm{i.i.d}}}{\sim}\left(X^{\prime j}-\nu_{-j}\left(X^{\prime-j}\right)\right)+\nu_{-j}(x^{-j})$, which is exactly the theoretical CPI sampling step, i.e. first we compute the conditional expectation $\nu_{-j}(x^{-j})$ with the regressor $\nu_{-j}$ and then we add a permuted residual $\left(X^{\prime j}-\nu_{-j}\left(X^{\prime-j}\right)\right)$. Then, for $\widetilde{X}^{(j)}\sim P^{\star}_{j}$, we have that $X^{j}\overset{\mathrm{i.i.d.}}{\sim}\widetilde{X}^{(j)}|X^{-j}$. We obviously have $X^{-j}=\widetilde{X}^{(j)-j}$. Finally, to ensure that $\widetilde{X}^{(j)}\perp\!\!\!\perp y\mid X^{-j}$, we can rely on the fact that, similar to the construction of knockoffs in Candès et al., (2018), $\widetilde{X}^{(j)}$ is constructed without using $y$.
@@ Appendix D Proofs of the double robustness
On the other hand, we observe that using Assumption 3.4 we have the equivalence between the conditional independence and functional independence, i.e. as $X^{j}\perp\!\!\!\perp y\mid X^{-j}$, then $m(X)$ does not depend on $X^{j}$. Let $\epsilon>0$. First, using the continuity of the loss, we have that for any $\epsilon$, there exists $\delta>0$ such that
@@ D.2 On the asymptotic relevance
In this section, we numerically explore the asymptotic relevance assumption (3.5). This assumption can be interpreted as requiring a model that does not extrapolate under the null hypothesis. Thanks to the equivalence between functional and conditional dependence (see Assumption 3.4 and Reyero-Lobo et al., (2025)), the theoretical model does not depend on the $j$-th coordinate whenever $X^{j}\perp\!\!\!\perp Y\mid X^{-j}$. This is easily illustrated for the linear model, where the coefficients of the null coordinates converge to zero; hence, imputations on these coordinates are irrelevant, since for sufficiently large $n_{0}$ the multiplied terms vanish.
In contrast, most models are not as transparent as the linear model, and one cannot simply verify the absence of dependence on the $j$-th coordinate by inspecting an estimated coefficient. However, this can be assessed model-agnostically using the Permutation Feature Importance (PFI, see Section B). The theoretical PFI index was proven to be zero under the null hypothesis in Reyero-Lobo et al., (2025), as it satisfies their proposed minimal axiom. In Figure 3, we show that for a large benchmark of ML models, the empirical PFI converges to zero. This indicates that there is no extrapolation under the null hypothesis, as the models indeed do not rely on the null coordinate.
@@
and we study the effect of varying the correlation between the null coordinate $X_{2}$ and the important coordinate $X_{1}$. In particular, we consider $\Sigma_{1,2}\in\{0,0.3,0.6,0.9\}$. The sample size is $n=1000$, with $n_{\mathrm{train}}=800$ and $n_{\mathrm{test}}=200$. All models are implemented using the default specifications of scikit-learn (Pedregosa et al., (2011)).
@@
We first note that using 3.7, we have that
@@ D.5.1 Linear decay of LOCO
Under 3.10 with $X$ Gaussian, we have that
@@
For an initial intuition, let’s assume we have independent training samples for each regressor (which is even proposed in Williamson et al., (2023) to control the type-I error).
@@
with $O_{P}(1/n)$ meaning that it decreases in probability linearly with $n$. We observe that in this example both errors follow the same distribution. This is because, as shown in Lemma I.2, the restricted model also satisfies the linear model. We observe that this is not generally the case: Lobo et al., (2024) showed that even under the global logistic assumption, the restricted model is not logistic.
@@ D.5.2 Quadratic decay of CPI
Appendix E Assumptions and proof of Theorem 4.2
In this section we present the assumptions taken from Williamson et al., (2023) to proof the asymptotic efficiency. To do so, we first introduce the same notations as in this paper.
@@
We also recall from Williamson et al., (2023) that Assumption A1 is referred to as optimality because we require a development that depends only on the second order, as the derivative tends to zero given that the model under consideration is a minimizer of the loss.
Similarly, as done in Williamson et al., (2023), we will prove the asymptotic efficiency of each empirical predictiveness measure with respect to its theoretical counterpart separately. This means that both the predictiveness measure using the information of the $j$-th covariate and the one without it will converge efficiently to their respective theoretical predictiveness measures. Therefore, all the previous assumptions need to be fulfilled for the restricted model. Additionally, we will require the minimum convergence rate not only of $\widehat{m}$, which was already required for the complete model convergence, but also for the model $\widehat{\nu}_{-j}$:
@@
We observe that all the above assumptions are exactly the ones required for the asymptotic efficiency of LOCO in Williamson et al., (2023), except for replacing the required minimum rate of convergence of $\widehat{m}_{-j}$ with that of $\widehat{\nu}_{-j}$.
@@
We recall that, in our setting—the knockoffs framework (Candès et al., (2018))—it is easier to fulfill the latter condition than the former. This is because the relationship between the input and the output is expected to be more complex than the relationship among the inputs.
For a detailed discussion of these assumptions, see Williamson et al., (2023).
@@
Similarly as done in Williamson et al., (2023), the asymptotic efficiency will be established by decomposing each predictiveness measure. Indeed, we are going to first prove that
@@
This comes directly from Theorem 2 of Williamson et al., (2023). Then, we need to proof this efficient asymptotic convergence
@@
Appendix G Inference: type-I error and power

### v5

@@ 6 Experiments
We study both simulated and real-data settings. In the former, the TSI can be computed explicitly (see Appendix K). All simulations are run over at least $50$ repetitions. The code is available at https://github.com/AngelReyero/Sobol-CPI.
All the proposed theory applies to any conditional sampler. To illustrate this, we implement Sobol-CPI with the regression sampler (3.1), which we denote by Sobol-CPI and with a Gaussian sampler from the fippy library: CFI (Strobl et al., 2008) and the conditional SAGE value function (scSAGEvf, Covert et al. (2021)). These correspond exactly to Sobol-CPI(1) and Sobol-CPI(50), respectively, but without the bias correction from Proposition 4.3, which we have added. Thus, they are equivalent to Sobol-CPI with a different sampler. We also compare to LOCO-W from Williamson et al. (2023), and LOCO-HD from Verdinelli and Wasserman (2024). Our focus is on bias in estimating important (efficiency) and non-important covariates (double robustness), as well as on enabling powerful, type-I-error-controlled variable selection. For a broader discussion on how $n_{\mathrm{cal}}$ balances nonparametric efficiency and double robustness, see Section J.2.
@@
in Figure 1 we study a nonlinear setting similar to Bénard et al. (2022): $y=X_{0}X_{1}\mathbb{I}_{X_{2}>0}+2X_{3}X_{4}\mathbb{I}_{X_{2}<0}$, with $X\sim\mathcal{N}(\mu,\Sigma)$, $\Sigma_{i,j}=0.6^{|i-j|}$, $p=50$, and $\mu=\mathbf{0}$. We use Lasso for $\widehat{\nu}_{-j}$ and a SuperLearner (van der Laan et al., 2007) of Random Forests, Lasso, Gradient Boosting, and SVM for $\widehat{m}$ and $\widehat{m}_{-j}$. For computational reasons, in the data-splitting LOCO-W, we use only Gradient Boosting. We apply Sobol-CPI with $n_{\mathrm{cal}}=1$ and $100$.
In this setting, since $X$ is Gaussian, Sobol-CPI with a Gaussian sampler (CFI and scSAGEvf) performs very well. Even using a complex model to estimate $\widehat{m}_{-j}$ but a simpler one for $\widehat{\nu}_{-j}$, Sobol-CPI not only achieves more computationally efficient estimates but also outperforms LOCO.
From the first plot, perturbation (Sobol-CPI(1)/CFI), removal (LOCO-W/HD), and marginalization (scSAGEvf/ Sobol-CPI(100)), converge to the TSI, justifying their comparison and the study of their statistical properties.
From the second plot, which highlights a null feature, and the fourth plot, where all null-feature importances are aggregated, the double robustness of Sobol-CPI is evident: refitting-based methods exhibit higher variability. This plot also shows PFI, which is omitted from the first plot since it does not estimate TSI. PFI correctly assigns no importance to the null features (second and fourth plots) and correctly identifies important features (third plot). These results shows that, even if the sampler is misspecified, the asymptotic relevance assumption (3.4) still holds.
Figure 1: Double robustness for complex learners: left to right: TSI estimates for an important covariate ($X_{0}$) and a null covariate ($X_{6}$); AUC for an importance-based variable selection; bias for null covariates.
@@
As discussed in Section 5, to test the null hypothesis, a correction is necessary to address the variance decrease under the null hypothesis. In Figure 2, we present the power with a linear correction term. Other corrections, along with analyses of type-I error, computation time, and additional settings (e.g., different correlations and polynomial settings), are discussed in Section J.4. Note that LOCO-W lacks power and requires a large sample size for type-I error control. In addition, Sobol-CPI also reduces the bias for the non-null covariates and leverages its double robustness property to achieve the largest power. Finally, the type-I error is controlled at the target level $0.05$ across all the procedures but the data-splitting version of Williamson et al. (2023).
Figure 2: Statistical Inference on variable importance in a linear setting: AUC for variable selection accuracy, bias in non-null TSI estimation, power and type-I error. Sobol-CPI(1) provides the most powerful test. Using the $O(1/n)$-corrected variance, the type-I error is controlled.
Real data:
we evaluate several standard ML models and datasets (see Section J.5.1); since conclusions were consistent, we report additional results in Appendix J.5. Here we present the Wisconsin Breast Cancer (WBDC) dataset with a SuperLearner. As the set of true important features is unknown in real data, we add an artificial null feature constructed as a noisy function of the original features, and vary its correlation with them (x-axis). This allows us to estimate null importance and type-I error on the artificial feature, while performing discoveries on the true inputs.
In Figure 3 we compare the $p$-value from LOCO-W, Sobol-CPI(1) with the Wilcoxon test, and the other procedures with a linear correction. This correction is conservative and yields no discoveries, whereas Sobol-CPI(1)-Wilcoxon SCPI(1)-wcx is powerful while still controlling the error. Also, observe that the double robustness holds in practice.
Figure 3: Double robustness and inference on real data: Using WBDC dataset with an added artificial null feature of varying correlation (x-axis). SCPI-based methods assign no importance to the null. With Wilcoxon test (Alg. 1), discoveries are made while controlling the error and remaining computationally efficient.
@@ 7 Conclusion
In this article, we revisited Conditional Feature Importance (CFI). We first studied the widely used regression-based estimator of the conditional distribution. We then analyzed both the null and alternative regimes. To explain the strong empirical performance of CFI in correctly identifying null features, we uncover a double robustness property: the method benefits jointly from the conditional sampler and the predictive model. Asymptotically, the sampler draws from the input distribution, so the two sources of error compensate by sharing the same distribution, while the model typically induces implicit variable selection during optimization.
Under the alternative regime, we modify CFI and introduce Sobol-CFI, clarifying its population target and proving nonparametric efficiency. We further address bias by proposing a correction that establishes a unified link between LOCO (refitting), CFI (perturbation), and SAGE value functions (marginalization). Finally, we show how to perform valid inference within this framework. Overall, CFI emerges as a powerful, statistically controlled variable importance measure, and this work significantly advances the understanding of the effect of conditional sampler estimation.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C Conditional sampling proofs
Also, we have that $X^{j}|\left(X^{-j}=x^{-j}\right)=\epsilon_{j}+\nu_{-j}(x^{-j})\overset{{\mathrm{i.i.d}}}{\sim}\left(X^{\prime j}-\nu_{-j}\left(X^{\prime-j}\right)\right)+\nu_{-j}(x^{-j})$, which is exactly the theoretical CPI sampling step, i.e. first we compute the conditional expectation $\nu_{-j}(x^{-j})$ with the regressor $\nu_{-j}$ and then we add a permuted residual $\left(X^{\prime j}-\nu_{-j}\left(X^{\prime-j}\right)\right)$. Then, for $\widetilde{X}^{(j)}\sim P^{\star}_{j}$, we have that $X^{j}\overset{\mathrm{i.i.d.}}{\sim}\widetilde{X}^{(j)}|X^{-j}$. We obviously have $X^{-j}=\widetilde{X}^{(j)-j}$. Finally, to ensure that $\widetilde{X}^{(j)}\perp\!\!\!\perp y\mid X^{-j}$, we can rely on the fact that, similar to the construction of knockoffs in Candès et al. (2018), $\widetilde{X}^{(j)}$ is constructed without using $y$.
@@ Appendix D Proofs of the double robustness
On the other hand, we observe that as $X^{j}\perp\!\!\!\perp y\mid X^{-j}$, then $m(X)$ does not depend on $X^{j}$. Let $\epsilon>0$. First, using the continuity of the loss, we have that for any $\epsilon$, there exists $\delta>0$ such that
@@ D.2 On the asymptotic relevance
In this section, we numerically explore the asymptotic relevance assumption (3.4). This assumption can be interpreted as requiring a model that does not extrapolate under the null hypothesis. The theoretical model does not depend on the $j$-th coordinate because $X^{j}\perp\!\!\!\perp Y\mid X^{-j}$. This is easily illustrated for the linear model, where the coefficients of the null coordinates converge to zero; hence, imputations on these coordinates are irrelevant, since for sufficiently large $n_{0}$ the multiplied terms vanish.
In contrast, most models are not as transparent as the linear model, and one cannot simply verify the absence of dependence on the $j$-th coordinate by inspecting an estimated coefficient. However, this can be assessed model-agnostically using the Permutation Feature Importance (PFI, see Section B). The theoretical PFI index was proven to be zero under the null hypothesis in Reyero-Lobo et al. (2025), as it satisfies their proposed minimal axiom. In Figure 4, we show that for a large benchmark of ML models, the empirical PFI converges to zero. This indicates that there is no extrapolation under the null hypothesis, as the models indeed do not rely on the null coordinate.
@@
and we study the effect of varying the correlation between the null coordinate $X_{2}$ and the important coordinate $X_{1}$. In particular, we consider $\Sigma_{1,2}\in\{0,0.3,0.6,0.9\}$. The sample size is $n=1000$, with $n_{\mathrm{train}}=800$ and $n_{\mathrm{test}}=200$. All models are implemented using the default specifications of scikit-learn ((Pedregosa et al., 2011)).
@@
In Figure 5 we fix the correlation level and observe that, as the sample size increases, all methods reduce their reliance on the null feature. This illustrates an asymptotic property: in practice, with finite samples and real data, we do not recommend using this sampler and instead advocate for conditional samplers. We further explore this assumption on real data in Appendix J.5.3.
Figure 5: Asymptotic relevance on standard ML models: Permutation Feature Importance (PFI) mean and standard deviation in parenthesis for $X_{1}$ (left) and $X_{2}$ (right) across different sample sizes and models, where $Y=X_{1}^{2}+\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.2),\quad X\sim\mathcal{N}(0,\Sigma),$ with $\Sigma_{1,2}=0.6$. The 80% of the sample size is used for training and 20% for testing. The experiment was repeated 100 times.
@@
We first note that using 3.6, we have that
@@ D.5.1 Linear decay of LOCO
Under 3.9 with $X$ Gaussian, we have that
@@
For an initial intuition, let’s assume we have independent training samples for each regressor (which is even proposed in Williamson et al. (2023) to control the type-I error).
@@
with $O_{P}(1/n)$ meaning that it decreases in probability linearly with $n$. We observe that in this example both errors follow the same distribution. This is because, as shown in Lemma I.2, the restricted model also satisfies the linear model. We observe that this is not generally the case: Lobo et al. (2024) showed that even under the global logistic assumption, the restricted model is not logistic.
@@ D.5.2 Quadratic decay of CPI
Appendix E Assumptions and proof of Theorem 4.2 (Nonparametric Efficiency)
In this section we present the assumptions taken from Williamson et al. (2023) to proof the asymptotic efficiency. To do so, we first introduce the same notations as in this paper.
@@
We also recall from Williamson et al. (2023) that Assumption A1 is referred to as optimality because we require a development that depends only on the second order, as the derivative tends to zero given that the model under consideration is a minimizer of the loss.
Similarly, as done in Williamson et al. (2023), we will prove the asymptotic efficiency of each empirical predictiveness measure with respect to its theoretical counterpart separately. This means that both the predictiveness measure using the information of the $j$-th covariate and the one without it will converge efficiently to their respective theoretical predictiveness measures. Therefore, all the previous assumptions need to be fulfilled for the restricted model. Additionally, we will require the minimum convergence rate not only of $\widehat{m}$, which was already required for the complete model convergence, but also for the model $\widehat{\nu}_{-j}$:
@@
We observe that all the above assumptions are exactly the ones required for the asymptotic efficiency of LOCO in Williamson et al. (2023), except for replacing the required minimum rate of convergence of $\widehat{m}_{-j}$ with that of $\widehat{\nu}_{-j}$.
@@
We recall that, in our setting—the knockoffs framework (Candès et al. (2018))—it is easier to fulfill the latter condition than the former. This is because the relationship between the input and the output is expected to be more complex than the relationship among the inputs.
For a detailed discussion of these assumptions, see Williamson et al. (2023).
@@
Similarly as done in Williamson et al. (2023), the asymptotic efficiency will be established by decomposing each predictiveness measure. Indeed, we are going to first prove that
@@
This comes directly from Theorem 2 of Williamson et al. (2023). Then, we need to proof this efficient asymptotic convergence
@@
Appendix G Inference
G.1 Variance correction: type-I error and power